OpenStudy (zairahernandez):

The environmental and health risks associated with ground-level ozone are greater than the environmental and health risks associated with the depletion of the ozone layer. True or false?

OpenStudy (abhisar):

False !

OpenStudy (anonymous):

i thought depletion of ozone layer caused the ground level ozone lol

OpenStudy (anonymous):

i probably sound stupid lol

OpenStudy (abhisar):

Ozone layer depletion causes increased exposure to UV rays which can cause Skin cancer or carcenoma Mutations While ground level ozone causes only breathing disorders.

OpenStudy (anonymous):

yup i sounded stupid lol

OpenStudy (abhisar):

no u are not \(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)